THE HIGH RESOLUTION JET COOLED SPECTRUM OF PENTAFLUOROETHANE

\underline{LISA M. GOSS}, Department of Chemistry, Idaho State University, Campus Box 8023, Pocatello ID, 83209; THOMAS A. BLAKE and ROBERT L. SAMS , Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, Richland, WA 99352.

The jet cooled spectrum of pentafluoroethane (C_{2}HF_{5}) has been recorded between 1100. and 1325. cm^{−1}at a resolution of 0.0022 cm^{−1}. A rotational temperature of approximately 10 K was achieved by expanding 50 Torr of C_{2}HF_{5} in 500 Torr of helium. Transitions belonging to five different vibrations have been assigned and fit to a Watson Hamiltonian: the ν_{3} at 1309.88 cm^{−1}, the ν_{4} at 1200.74 cm^{−1}, the ν_{5} at 1142.78 cm^{−1}, the ν_{13} at 1223.33 cm^{−1}, and the ν_{14} at 1147.39 cm^{−1}. The rms deviations of the fits of the ν_{4} band (0.0004 cm^{−1}) as well as the overlapping ν_{5} (0.0006 cm^{−1}) and ν_{14} (0.0004 cm^{−1}) bands are in satisfactory agreement with the experimental uncertainty. The perturbations observed by McNaughton et al^{a} are still apparent in the ν_{3} and ν_{13} at the rotational temperature of this study but satisfactory fits have been achieved.

aD. McNaughton et al., Vib. Spec., \underline{36}, 123 (2004); C. D. Thompson et al., J. Mol. Spec., \underline{230}, 133 (2005).